Gross National Expenditure (GNE) 1140, Depreciation 100, Direct taxes 100, Indirect taxes 54,Total Imports 200, Transfer payments 200, Total Exports 100 - What would be the GDP??
Assuming there is investment in the preceding years of 40, the multiplicative effects of investment result in total output of 2364 of which 100 is kept aside for future years inventory & 200 as intermediate inputs. Raw material inputs to industries was 300 & the corporate losses are 200. Indirect taxes are 186 & wages are 500. How would this impact GDP??
1
Expert's answer
2013-04-15T09:37:27-0400
Personal savings (S) plus personal consumption (C) = personal disposable income (PDI)PDI plus personal taxes paid minus transfer payments received = personal income (PI)PI plus retained earnings plus corporate taxes plus transfer payments plus interest on the public debt = net national income (NNI)NNI plus indirect taxes = net national product (NNP)NNP plus depreciation = gross national product (GNP)
So, GDP = GNE + Depreciation + Taxes + Transfer payments + (E - I) = 1140 +100 + 154 + 200 + (100 - 200) = 1494

In the second case GDP = 40 + (2364 - 100 - 200) - 200 + 186 + 500 = 2590

Raw materials are not included in the calculation.

So, GDP increased after the investment.